# Optimal control of the fidelity coefficient in a Cahn–Hilliard image inpainting model

### v1

@@ 4.3 Adjoint system
In this way, we can get the equations above. We also mention that if $\alpha_{2}>0$, due to the low regularity of the final datum $p(T)$, which can be at most in $L^{\infty}(\Omega)$, we are only able to establish the existence of very weak solutions to (4.36)–(4.39). This is enough to derive first-order necessary conditions. However, we will need higher regularity for second-order ones. For this reason, we will also assume $\alpha_{2}=0$ in the following. One could keep $\alpha_{2}>0$ if, for some particular application, the target image $f$ could be replaced by a regularised version $\tilde{f}\in H^{2}(\Omega\setminus D)$, for instance through the system (A.1)–(A.3). However, it is not clear if this could be beneficial in general, therefore we stick to the standard case $f\in L^{\infty}(\Omega\setminus D)$. Finally, note that most classical PDE-based inpainting methods impose either a boundary condition on $\Omega\setminus D$ or a penalty term that keeps $\varphi$ close to $f$ for all time. Hence, many of the references in the literature can be viewed as the special case $\alpha_{2}=0$.
@@
Let A1A3, A6 and C1C2 hold Consider the strong solution $(\overline{\varphi},\overline{\mu})\in\mathbb{X}$ to (2.5)–(2.8) corresponding to $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$. Then the adjoint problem (4.36)–(4.39) admits a unique weak solution such that
@@
  • If $g_{1}\in L^{2}(0,T;H)$ and $g_{2}\in W$, then there exists a unique strong solution to (4.50)–(4.53) such that
@@ 4.4 First-order optimality conditions
$\int_{0}^{T}\int_{\Omega\setminus D}\alpha_{1}p(f-\overline{\varphi})(\lambda-\overline{\lambda}_{0})\,\mathrm{d}x\,\mathrm{d}t-\int_{\Omega\setminus D}\frac{\beta}{\overline{\lambda}_{0}^{3}}(\lambda_{0}-\overline{\lambda}_{0})\,\mathrm{d}x\geq 0\quad\hbox{for any $\lambda_{0}\in\mathcal{U}_{\text{ad}}$,}$ (4.56)
@@ 5 Second-order optimality conditions
Observe that problem (2.5)–(2.8) is nonlinear and so is the control-to-state operator $\mathcal{S}$. Therefore, our optimal control problem (CP) is not convex and the first-order conditions given by Theorem 4.8 are only necessary and not sufficient. In this section, we derive second-order sufficient conditions to characterise the optimal solutions of (CP).
5.1 Control-to-costate operator
In view of computing the second-order derivatives of the cost functional $\mathcal{J}$, we now introduce the control-to-costate operator. Since we require more regularity (see Remark 4.5), from now on we assume that
@@ 5.1 Control-to-costate operator
Then, we define the control-to-costate operator by setting, for any $\lambda_{0}\in\mathcal{U}_{\text{ad}}$,
@@
$\quad+C\lVert\varphi_{1}-\varphi_{2}\rVert^{2}_{L^{\infty}(0,T;V)}\int_{0}^{T}\lVert q_{2}\rVert^{2}_{V}\,\mathrm{d}t$
@@
so that (5.19) holds. and the proof is complete. ∎
@@ 5.2 Second-order optimality conditions
We are now ready to state and prove the second-order sufficient optimality conditions for our optimal control problem (CP). To do this, we follow the approach developed in [8, 9] (cf. also [34] for a slightly different one). Both were later applied also to Cahn–Hilliard-type systems (see [1, 18, 16]). The idea of sufficient second-order conditions is to show that if $\mathrm{D}^{2}\mathcal{J}_{\text{red}}(\overline{\lambda}_{0})$ is positive definite in an optimal point $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$, namely
@@
then such a $\overline{\lambda}_{0}$ is a strict local minimiser of $\mathcal{J}_{\text{red}}$ on $\mathcal{U}_{\text{ad}}$. However, imposing the positive definiteness of $\mathrm{D}^{2}\mathcal{J}_{\text{red}}$ in all directions $h_{0}$ is generally too restrictive. Indeed, it suffices to require such a condition only in a certain class of critical directions (cf. [34, Section 4.10] and [8, 9] for more details on this topic). Essentially, due to some additional information on the structure of the optimal controls that can be extracted from the first-order optimality condition and from the fact that $\mathcal{U}_{\text{ad}}$ has a box form (cf. [34, 8, 9]), one can restrict to impose the positive definiteness only in some particular directions. Hence, we need to introduce some notions which are summarized in the following
@@
Then, by well-known compact embeddings (cf. [32, Section 8, Corollary 4]), it follows that, up to a subsequence,
@@ 6 Conclusions
In this paper, we introduced and studied an optimal control approach to the image inpainting model proposed by Bertozzi, Esedoglu, and Gillette. Our primary focus was on determining a suitable fidelity coefficient $\lambda$ to achieve faithful reconstruction of a damaged image $f$. By moving from the original polynomial double-well potential to a physically relevant logarithmic-type potential, we ensured that solutions to our evolutionary system remained strictly within the interval $[-1,1]$, thus preserving the range of the inpainted image at all times. After establishing the well-posedness of the state system, we formulated an optimal control problem (CP) whose goal was twofold: maintain fidelity to the undamaged image $f$ in $\Omega\backslash D$ and promote sufficiently large values of $\lambda$ where fine details of the image need to be restored. We proved the existence of an optimal control $\lambda$ under the proposed constraints and then derived first-order necessary optimality conditions using an adjoint system. Subsequently, we addressed second-order sufficient conditions, which provide deeper insight into the local optimality of solutions.
@@ Appendix A Appendix
Acknowledgments. C. Cavaterra, M. Fornoni, and M. Grasselli are members of GNAMPA (Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni) of INdAM (Istituto Nazionale di Alta Matematica). C. Cavaterra, M. Fornoni and M. Grasselli have been partially supported by the MIUR-PRIN Grant 2020F3NCPX “Mathematics for Industry 4.0 (Math4I4)”. C. Cavaterra has been partially supported by the MIUR-PRIN Grant 2022 “Partial differential equations and related geometric-functional inequalities”. Elena Beretta’s research has been partially supported by NYUAD Science Program Project Fund AD364.
This research is part of the activities of “Dipartimento di Eccellenza 2023-2027” of Università degli Studi di Milano (C. Cavaterra) and Politecnico di Milano (M. Grasselli).

### v2

@@ 4.3 Adjoint system
In this way, we can get the equations above. We also mention that if $\alpha_{2}>0$, due to the low regularity of the final datum $p(T)$, which can be at most in $L^{\infty}(\Omega)$, we are only able to establish the existence of very weak solutions to (4.36)–(4.39). This is enough to derive first-order necessary conditions. However, we will need higher regularity for second-order ones. For this reason, we will also assume $\alpha_{2}=0$ in the following (see Section 5).
@@
Let A1A3, A6 and C1C2 hold. Consider the strong solution $(\overline{\varphi},\overline{\mu})\in\mathbb{X}$ to (2.5)–(2.8) corresponding to $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$. Then the adjoint problem (4.36)–(4.39) admits a unique weak solution such that
@@
  • If $g_{2}\in L^{2}(0,T;H)$ and $g_{4}\in W$, then there exists a unique strong solution to (4.50)–(4.53) such that
@@ 4.4 First-order optimality conditions
$\int_{0}^{T}\int_{\Omega\setminus D}\alpha_{1}p(f-\overline{\varphi})(\lambda_{0}-\overline{\lambda}_{0})\,\mathrm{d}x\,\mathrm{d}t-\int_{\Omega\setminus D}\frac{\beta}{\overline{\lambda}_{0}^{3}}(\lambda_{0}-\overline{\lambda}_{0})\,\mathrm{d}x\geq 0\quad\hbox{for any $\lambda_{0}\in\mathcal{U}_{\text{ad}}$,}$ (4.56)
@@ 5 Second-order optimality conditions
Observe that problem (2.5)–(2.8) is nonlinear and so is the control-to-state operator $\mathcal{S}$. Therefore, our optimal control problem (CP) is not convex and the first-order conditions given by Theorem 4.8 are only necessary and not sufficient. In this section, we derive second-order sufficient conditions to characterise the optimal solutions of (CP). In general, such conditions require more regularity on the solutions to the adjoint system (4.36)–(4.39). In view of Remark 4.7, we observe that one could keep $\alpha_{2}>0$ if, for some particular application, the target image $f$ could be replaced by a regularised version $\tilde{f}\in H^{2}(\Omega\setminus D)$, for instance through the system (A.1)–(A.3). However, it is not clear if this could be relevant from the point of view of applications (e. g. numerical approximations), therefore we stick to the standard case $f\in L^{\infty}(\Omega\setminus D)$. This choice forces us to assume that
@@
Note that most classical PDE-based inpainting methods impose either a boundary condition on $\Omega\setminus D$ or a penalty term that keeps $\varphi$ close to $f$ for all time. Hence, many of the references in the literature can be viewed as the special case $\alpha_{2}=0$.
5.1 Control-to-costate operator
In view of computing the second-order derivatives of the cost functional $\mathcal{J}$, we now introduce the control-to-costate operator. Then, for any $\lambda_{0}\in\mathcal{U}_{\text{ad}}$, we define
@@ 5.1 Control-to-costate operator
$\quad+C\lVert\overline{\varphi}_{1}-\overline{\varphi}_{2}\rVert^{2}_{L^{\infty}(0,T;V)}\int_{0}^{T}\lVert q_{2}\rVert^{2}_{V}\,\mathrm{d}t$
@@
so that (5.19) holds and the proof is complete. ∎
@@ 5.2 Second-order optimality conditions
We are now ready to state and prove the second-order sufficient optimality conditions for our optimal control problem (CP). To do this, we follow the approach developed in [8, 9] (cf. also [11, 38] for a slightly different one). Both were later applied also to Cahn–Hilliard-type systems (see [1, 21, 18, 36, 16, 19]). The idea of sufficient second-order conditions is to show that if $\mathrm{D}^{2}\mathcal{J}_{\text{red}}(\overline{\lambda}_{0})$ is positive definite in an optimal point $\overline{\lambda}_{0}\in\mathcal{U}_{\text{ad}}$, namely
@@
then such a $\overline{\lambda}_{0}$ is a strict local minimiser of $\mathcal{J}_{\text{red}}$ on $\mathcal{U}_{\text{ad}}$. However, imposing the positive definiteness of $\mathrm{D}^{2}\mathcal{J}_{\text{red}}$ in all directions $h_{0}$ is generally too restrictive. Indeed, it suffices to require such a condition only in a certain class of critical directions (cf. [38, Section 4.10] and [8, 9] for more details on this topic). Essentially, due to some additional information on the structure of the optimal controls that can be extracted from the first-order optimality condition and from the fact that $\mathcal{U}_{\text{ad}}$ has a box form (cf. [38, 8, 9]), one can restrict to impose the positive definiteness only in some particular directions. Hence, we need to introduce some notions which are summarized in the following
@@
Then, by well-known compact embeddings (cf. [35, Section 8, Corollary 4]), it follows that, up to a subsequence,
@@ 6 Conclusions
In this paper, we introduced and studied an optimal control approach to the image inpainting model proposed by Bertozzi, Esedoglu, and Gillette. Our primary focus was on determining a suitable fidelity coefficient $\lambda$ to achieve faithful reconstruction of a damaged image $f$. By moving from the original polynomial double-well potential to a physically relevant logarithmic-type potential, we ensured that solutions to our evolutionary system remained strictly within the interval $[-1,1]$, thus preserving the range of the inpainted image at all times. After establishing the well-posedness of the state system, we formulated an optimal control problem whose goal was twofold: maintain fidelity to the undamaged image $f$ in $\Omega\backslash D$ and promote sufficiently large values of $\lambda$ where fine details of the image need to be restored. We proved the existence of an optimal control $\lambda$ under the proposed constraints and then derived first-order necessary optimality conditions using an adjoint system. Subsequently, we addressed second-order sufficient conditions, which provide deeper insight into the local optimality of solutions.
@@ Appendix A Appendix
Acknowledgments. The authors wish to thank the anonymous reviewers, who carefully read the manuscript and provided many comments that improved the quality of the paper. C. Cavaterra, M. Fornoni, and M. Grasselli are members of GNAMPA (Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni) of INdAM (Istituto Nazionale di Alta Matematica). C. Cavaterra, M. Fornoni and M. Grasselli have been partially supported by the MIUR-PRIN Grant 2020F3NCPX “Mathematics for Industry 4.0 (Math4I4)”. C. Cavaterra has been partially supported by the MIUR-PRIN Grant 2022 “Partial differential equations and related geometric-functional inequalities”. Elena Beretta’s research has been partially supported by NYUAD Science Program Project Fund AD364. This research is part of the activities of “Dipartimento di Eccellenza 2023-2027” of Università degli Studi di Milano (C. Cavaterra) and Politecnico di Milano (M. Grasselli).